OpenStudy (jessie13):

Will Medal Answer ASAP!!!!! please help! 1: What is the GCF of the numerator and denominator in the following fraction? -99/176 3 9 11 They are relatively prime. 2: Reduce -56/-70 to lowest terms. 8/10 -4/5 -2/3 4/5 3: Reduce -34/87 to lowest terms. -2/5 -1/3 -11/29 It is already reduced.

OpenStudy (jessie13):

@welshfella

OpenStudy (welshfella):

factor -99 and 176 -99 = -3*3*11 176 = 2*2*2*11 to find the GCF pick out common factors

OpenStudy (jessie13):

11

OpenStudy (welshfella):

what is in the factors of both -99 and 176?

OpenStudy (welshfella):

right that's it :- 11 is the GCF

OpenStudy (jessie13):

okay thanks do u know #2 and #3?

OpenStudy (welshfella):

for number 2 find the factors of 56 and 70 and cancel out the common ones also remember that - / - = +

OpenStudy (jessie13):

??????

OpenStudy (welshfella):

so we have